# Aptly: Making Mobile Apps from Natural Language

### v1

@@ III-B2 Scenario-Based Task
Participants provided initial descriptions that ranged from simple one-liners like “Make an app that is a calculator” to detailed paragraphs specifying the layout and functionality, including voice recognition and auditory feedback. This task demonstrated the ability of participants to conceptualize and articulate diverse solutions using Aptly, catering to a specific accessibility need. Participants’ approaches varied; some chose to incrementally build and refine their app’s user interface and functionalities using Aptly’s editing tools, while others outlined their complete vision from the onset. For example, one participant started with a basic calculator design and progressively enhanced its features and interface to be more accessible for visually impaired users. This approach highlights Aptly’s flexibility to accommodate different development styles and the iterative nature of designing user-centric applications.
Moreover, the participants’ creative solutions for accessibility features showcased the potential of integrating advanced technologies such as text-to-speech and tactile feedback into apps. One innovative approach involved using screen taps and phone vibrations to communicate numbers and results, demonstrating how users can leverage Aptly not only as a development tool but also as a platform for ideation. The diversity in usage and the degree of reliance on Aptly’s capabilities varied, indicating that while some users fully embraced the AI-driven development process, others preferred a more hands-on approach in shaping their applications. This underscores the importance of versatile tools like Aptly in modern app development, especially when addressing the specific needs of diverse users.
III-B3 General Observation
Observations revealed valuable insights into user behavior and platform design. Participants exhibited a general impatience with the system’s response times, often proceeding to plan subsequent changes if immediate implementation of edits was not observed. Challenges arose when Aptly misunderstood user commands; while some participants attempted to rephrase their inputs multiple times using Aptly editing, others opted to manually make the changes. Notably, more experienced programmers frequently preferred manual adjustments and issued a higher number of commands compared to their less experienced counterparts. These seasoned programmers expressed a desire for greater control, including access to error messages to better understand the system’s handling of their commands. This feedback highlights a tension between designing intuitive interfaces suitable for beginners and providing detailed control and feedback that experienced users seek. This dichotomy suggests that while simplifying interfaces can benefit novices, it may frustrate advanced users who expect deeper interaction and transparency from the development tools.
III-B4 Participant responses
@@
  • “Programming is accessible for everyone”: Prior to the workshop, students’ views on programming accessibility varied, with some expressing skepticism about universal accessibility due to factors such as technological availability and socioeconomic barriers. Common responses ranged from “somewhat disagree” to acknowledging partial agreement, contingent on access to technology like computers and the internet. After using Aptly, participants generally showed a nuanced view, with some noting improvements in their perceptions of accessibility. Notably, though some recognized improvements, they still highlighted limitations, particularly for individuals who might lack basic digital literacy or access to necessary hardware.
@@
  • “Artificial Intelligence can be a powerful tool to help people in programming”: Initially, responses indicated a positive sentiment towards AI as a powerful aid in programming, with a slight variety hinting at potential job displacement concerns. After the workshop, reflections were more detailed, emphasizing the practical benefits and challenges observed during the use of AI-driven tools like Aptly. Participants appreciated AI’s potential to simplify the programming process, especially for beginners, yet also discussed the importance of understanding underlying logic beyond AI assistance. This shift suggests a more informed appreciation of AI’s capabilities and limitations in programming contexts.
We also asked each participant the question “Did Aptly change your interest in learning how to program?” on a scale of 1 to 5, where higher values signify increased interest. The range of responses spanned from a low of 2 to a high of 4, with no extreme sentiments expressed, as the ratings did not include the lowest (1) or highest (5) possible marks. This distribution implies that while Aptly was generally effective in increasing interest among most participants, it had a varied impact. The mean response was 3.7 (SD = 0.67). The mixed response highlights the potential of Aptly to engage users, albeit not uniformly, suggesting avenues for further investigation into individual differences and tool optimization to broaden its appeal and effectiveness.
III-B5 Limitation of Aptly based on responses
In our study, participants identified several limitations of the Aptly platform, providing valuable insights into its operational shortcomings and areas ripe for development. Notably, Aptly struggled with complex command executions such as playing specific audio files, drawing predefined shapes, and making API calls. Furthermore, it demonstrated limitations in processing abstract or intricate commands, which underscores deficiencies in its natural language processing capabilities.
Feedback from users on improving Aptly emphasized enhancing both functionality and user interface design. Participants suggested increasing the accuracy and range of command interpretation to reduce errors, thereby making the platform more robust for complex programming tasks. Improvements in the user interface were recommended to make the tool more intuitive and accessible, with clearer design elements that could facilitate easier navigation and usage. Enhanced error messaging and user feedback, particularly for users with programming experience, were noted as critical for allowing more effective debugging and application refinement.
These findings indicate a significant need for advancements in natural language understanding within Aptly to accommodate a broader array of user instructions and technical demands. The study highlights the essential balance between developing a tool that is accessible for beginners while being sufficiently advanced to support more complex user requirements. Expanding functionalities to include sophisticated web operations and comprehensive multimedia management, along with providing educational examples or templates, could help users better understand and utilize Aptly’s full range of capabilities. This comprehensive feedback not only reflects the current usability challenges faced by Aptly but also provides a strategic roadmap for future enhancements to better serve a diverse user base for tool developers.
IV Related Work
Automatic programming [17, 18], with its rich historical tapestry, has always been about bridging the communication gap between human intent and machine execution. At its core, it embodies the principle of transforming higher-level instructions into forms that a computer can readily understand and act upon. Today’s automatic programming aims to streamline the software development process, enabling developers to emphasize problem definition over intricate implementation details. This evolution promises a future of enhanced collaboration between humans and machines in software creation, making the process more accessible, efficient, and resilient against errors.
Modern automatic programming systems leverage machine learning models to generate code from a higher-level specification or even from natural language descriptions [19]. This has broadened the scope of automatic programming from simple code translation to complex code synthesis, opening up new possibilities for software development [20, 21]. Take, for instance, GitHub Copilot, a coding assistant powered by OpenAI Codex [22]. GitHub Copilot offers contextually relevant code suggestions as you type, effectively automating part of the coding process. Whether you’re learning a new language or building a complex system, Copilot serves as a knowledgeable companion, aiding in writing new code, navigating existing code, and even generating unit tests [23, 24, 25, 26, 27, 28, 29, 30].
Aptly is an attempt to push such boundaries to automatically create block code based on the user’s utterance. Text2App is a similar system based on natural language processing to convert user input into a representation eventually compatible with MIT App Inventor [31]. Such tools are moving toward a direction where anyone can create functional software without any need for traditional coding [32].
V Discussion & Conclusion
Overall, the design and implementation of tools like Aptly should inform how LLMs may apply to other visual programming environments, whether they are block based, flow based, or some other format. Our findings are most applicable to other block based programming environments built on Google Blockly, but the concepts should be translatable to other visual paradigms as well. By designing Aptly to be like Python while maintaining a correspondence to the original visual environment, we are able to leverage the depth of knowledge of LLMs to speed system development. We were also surprised by how well the LLMs were able to transform non-English instructions into working code examples as none of our example prompts were expressed in any language other than English. This suggests that we may see a much lower barriers to entry in the future with regard to computer science education as LLMs gain adoption.
In this work, we described the Aptly language, architecture, and server, and how they form a single system alongside the MIT App Inventor visual programming environment. We demonstrated how large language models are able to synthesize programs in this new language without any training on a new corpus. We further showed how edits to Aptly programs can be translated into corresponding edits in App Inventor projects, enabling Aptly to act as a collaborative programming agent. A small pilot study showed that students were better able to articulate how they would leverage AI systems like Aptly after working with the system, but that there was not a significant shift in their overall interest in learning to program.
The advent of no-code platforms such as Aptly poses both opportunities and challenges for computer science education, which merit careful exploration and discussion. As this trend continues to revolutionize the software industry, it’s becoming increasingly clear that our approach to computer science education may need to evolve to keep pace with these changes [33]. Traditional computer science education has long been centered on manual coding and algorithm design. It emphasizes the need for students to understand the intricacies of programming languages, the efficiency of algorithms, and the design of complex software systems. AI-based code-generation techniques may incentivize students to bypass such productive thinking. Future research on students’ usage of such tools, what they learn while using them [34], and how they feel about using these tools [35] will be active research.
Acknowledgment
We thank the whole MIT App Inventor team who helped throughout the development process, especially our staff and students including Evie Cai, Jacky Chen, Mark Friedman, Maura Kelleher, Jasmine Shone, Joyce Yuan. Thank you also to OpenAI and Drew Houston for their support and resources.
The second listing of Section II-C was generated using Aptly’s generative AI capabilities.

### v2

@@ 4.2. Scenario-Based Task
Participants provided initial descriptions that ranged from simple one-liners like “Make an app that is a calculator” then incrementally build and refine their app’s user interface and functionalities using Aptly’s editing tools to detailed paragraphs specifying the layout and functionality, including voice recognition and auditory feedback (e.g., “Have a third button that hears one number, then hears either add or subtract, then hears a second number, after all three components are heard, then display the sum or difference of the two numbers in the top, then display the number in the display box too”).
An interesting observation is that the participants’ creative solutions for accessibility features were varied. While most of the participants added audio systems for the accessibility feature, one innovative approach involved using screen taps and phone vibrations to communicate numbers and results (“Make each displayed element bigger. Also, when the result is displayed, vibrate the number of times of the result”).
5. Dicussion
Observations revealed valuable insights into user behavior and platform design. Some participants exhibited impatience with the system response times as it generally takes around $3\sim 10$ seconds for a change to reflect. They often proceed to prompt subsequent changes if an immediate implementation of the edits was not observed. Challenges arose when Aptly misunderstood user commands; while some participants attempted to rephrase their inputs multiple times using Aptly editing, others opted to manually make the changes. Notably, more experienced programmers frequently preferred manual adjustments and issued a higher number of commands compared to their less experienced counterparts. These seasoned programmers expressed a desire for greater control, including access to error messages to better understand the system’s handling of their commands. This feedback highlights a tension between designing intuitive interfaces suitable for beginners and providing detailed control and feedback that experienced users seek. This dichotomy suggests that while simplifying interfaces can benefit novices, it may frustrate advanced users who expect deeper interaction and transparency from the development tools.
5.1. Participant responses
@@
  • “Programming is accessible for everyone”: Prior to the workshop, students’ views on programming accessibility varied, with some expressing skepticism about universal accessibility due to factors such as technological availability and socioeconomic barriers. Common responses ranged from “somewhat disagree” to acknowledging partial agreement. After using Aptly, participants generally showed a nuanced view, with some noting improvements in their perceptions of accessibility. Notably, though some recognized improvements, they still highlighted limitations, particularly for individuals who might lack basic digital literacy or access to necessary hardware.
@@
  • “Artificial Intelligence can be a powerful tool to help people in programming”: Initially, responses indicated a positive sentiment towards AI as a powerful aid in programming, with a slight variety hinting at potential job displacement concerns. After the experiment, reflections were more detailed, emphasizing the practical benefits and challenges observed during the use of AI-driven tools like Aptly. Participants appreciated AI’s potential to simplify the programming process, especially for beginners, yet also discussed the importance of understanding underlying logic beyond AI assistance.
5.2. Limitation of Aptly based on responses
In our study, participants identified several limitations of the Aptly platform, providing valuable insights into its operational shortcomings and areas ripe for development. Notably, Aptly struggled with complex command executions such as playing specific audio files, drawing predefined shapes, and making API calls. Improvements in the user interface were recommended to make the tool more intuitive. Especially users were confused whether their prompt was not clear enough or the system is not capable to conduct their request. Enhanced error messaging and user feedback, particularly for users with programming experience, were noted as critical for allowing more effective debugging and application refinement. These findings indicate a significant need for advancements in natural language understanding within Aptly to accommodate a broader array of user instructions and technical demands.
6. Conclusion
This research introduces Aptly, a platform designed to empower individuals, particularly young learners, to create mobile applications that positively impact their lives and communities. By integrating a Large Language Model (LLM) with the App Inventor platform, Aptly simplifies the app development process. This involved engineering a novel textual language that directly corresponds with App Inventor’s block-based programming and employing effective prompt engineering to teach the model the platform’s functionality.
A preliminary study with ten high school students revealed valuable insights into Aptly’s usability and potential. Notably, programming experience influenced how students interacted with the platform. Participants identified areas for improvement, such as enhanced error handling and expanded capabilities like audio file integration. Significantly, Aptly shifted many participants’ perspectives on both programming accessibility and the role of AI in this process. This study provides crucial guidance for Aptly’s continued development and lays the groundwork for future research exploring its potential to facilitate computational action and empower a wider range of individuals to engage in app creation.